# Semi-supervised Graph Anomaly Detection via Robust Homophily Learning

### v1

@@ Abstract
Semi-supervised graph anomaly detection (GAD) utilizes a small set of labeled normal nodes to identify abnormal nodes from a large set of unlabeled nodes in a graph. Current methods in this line posit that 1) normal nodes share a similar level of homophily and 2) the labeled normal nodes can well represent the homophily patterns in the normal class. However, this assumption often does not hold well since normal nodes in a graph can exhibit diverse homophily in real-world GAD datasets. In this paper, we propose RHO, namely Robust Homophily Learning, to adaptively learn such homophily patterns. RHO consists of two novel modules, adaptive frequency response filters (AdaFreq) and graph normality alignment (GNA). AdaFreq learns a set of adaptive spectral filters that capture different frequency components of the labeled normal nodes with varying homophily in the channel-wise and cross-channel views of node attributes. GNA is introduced to enforce consistency between the channel-wise and cross-channel homophily representations to robustify the normality learned by the filters in the two views. Experiments on eight real-world GAD datasets show that RHO can effectively learn varying, often under-represented, homophily in the small normal node set and substantially outperforms state-of-the-art competing methods. Code is available at https://github.com/mala-lab/RHO.
@@ 1 Introduction
Graph anomaly detection (GAD) has received increasing attention in recent years due to its wide range of applications, such as fraud detection in finance, review spam detection, and abusive user detection [1, 26, 19, 35]. Semi-supervised GAD, which aims to leverage a small set of labeled normal nodes to identify abnormal nodes from a large set of unlabeled nodes in a graph, is among the most realistic problem settings in real-life applications [32, 36, 35]. This is because GAD datasets are typically dominated by normal nodes, which can greatly facilitate the annotation of a few normal nodes, e.g., if we randomly sample a few nodes from a large graph, most nodes would be normal. Current semi-supervised GAD methods are generally built upon two key assumptions: (1) all normal nodes exhibit a similar level of homophily, and (2) the partially labeled normal nodes can well represent the overall homophily pattern of the entire graph [11, 6, 36, 5, 40]. However, these two assumptions often do not hold well in the real-world GAD datasets. This is because although normal nodes generally exhibit high homophily, there can be large variations of homophily across the normal nodes, some of which can possess relatively low homophily (see Fig. 1a for visualization of this observation on the datasets, Amazon [9] and Elliptic [41]. More visualization results can be found in App. D). As a result, the GAD methods that rely on the aforementioned assumptions learn inaccurate homophily patterns of the normal class, leading to the misclassificaiton of low-homophily normal nodes as abnormal. This is particularly true when they use the popular neighborhood aggregation mechanisms that are prone to produce oversmooth homophily representations [35].
@@
Figure 1: (a) Homophily distribution of normal nodes in Amazon [9] and Elliptic [41]. (b) and (c) AUC comparison of GCN filter [40], BWGNN filter [39], and AdaFreq filter on the two datasets under three cases of homophily levels. The three cases are normal nodes sampled from the sets of low- and high homophily normal nodes, where nodes with homophily greater than 0.9 for Amazon and 0.7 for Elliptic are considered as high homophily. The datasets in the three cases are sampled using a ratio of (80%, 20%), (50%, 50%), and (20%, 80%) to the two node sets, respectively.
RHO consists of two novel modules, including adaptive freqency response filters (AdaFreq) and graph normality alignment (GNA). Unlike existing filters that are bounded by prior knowledge of a specific homophily, AdaFreq learns a set of adaptive filters with learnable parameters to robustly fit normal nodes of varying levels of homophily. For example, as illustrated in Fig. 1b and 1c, conventional GCN filters [40] rely on a low-frequency assumption of normal nodes, while the spectral filters in BWGNN [39] learns a pre-defined type of frequency response, both of which fail to learn generalized representations for normal nodes across three different homophily variations; AdaFreq is instead optimized with learnable parameters to fit adaptively to the normal nodes with three different homophily cases, showing much more robustness than the other two methods. In RHO, AdaFreq is utilized to learn such adaptive filters in both channel-wise and cross-channel views of node attributes, enabling the learning of heterogeneous normal patterns in diverse feature channels of the labeled normal nodes.
The set of filters learned in AdaFreq can differ from each other substantially. To obtain robust yet consistent representations of the normal nodes, the GNA module is introduced to enforce consistency of homophily representations learned across the filters. To this end, GNA constructs positive pairs using the representations from the channel-wise and cross-channel views at the corresponding nodes, and then maximizes the similarity between the representations learned from these two views while minimizing the similarity of negative pairs composed by representations from different nodes.
In doing so, RHO learns heterogeneous normal representations of normal nodes with varying degrees of homophily via AdaFreq, while ensuring the consistency of the learned normality via GNA. In summary, this work makes the following main contributions.
@@ 2.1 Graph Anomaly Detection (GAD)
Non-spectral GAD Methods. Many non-sepctral GAD methods apply traditional anomaly detection techniques on a GNN backbone, such as reconstruction [6, 11], adversarial learning [5, 4], one-class classification [46, 40, 2]. Other methods are based on measures designed for GAD, such as the recently proposed affinity-based methods [31, 34, 30, 29]. These methods are typically designed for the unsupervised setting, where no labeled nodes are available. They can be easily adapted to the semi-supervised setting by refining their objective to the labeled normal nodes. GGAD [36] is the first method specifically designed for the semi-supervised setting by generating the outliers and training a discriminative one-class classifier with the given labeled normal nodes. These methods failed to consider the homophily discrepancy among the normal nodes. Our proposed RHO addresses this by employing an adaptive filter that learns normal representations from both cross-channel and channel-wise views through a one-class optimization objective.
Spectral GAD Methods. The spectral GAD methods focus on the filter design in GNNs to have a more effective learning the representations of nodes [3, 38, 12]. Among them AMNet [3], BWGNN [38], and GHRN [12] aim to utilize the frequency signal using different spectral-based GNNs to distinguish between normal and abnormal nodes. Although these spectral-based GAD methods have achieved remarkable success [23, 8], they are typically designed for fully supervised setting where both labeled normal and abnormal nodes are available during training. On the other hand, the filters they utilize learn only predefined frequency information, making them difficult to learn expressive representations for normal nodes of different frequency components. RHO leverages AdaFreq to assign learnable parameters on different feature channels to effectively learn heterogeneous normal representations from the set of normal nodes of varying homophily.
@@ 2.2 Graph Homophily Modeling for GAD
Graph homophily modeling methods aim to address the homophily discrepancy in graphs, where target nodes often connect to nodes from different classes [43, 13, 25]. This inter-class connectivity negatively affects the quality of node representations within each class. The problem becomes more pronounced in GAD, where the homophily gap is typically large due to the inherent imbalance in GAD datasets [43, 24, 35]. Existing studies on Graph homophily modeling for GAD primarily focus on neighbor selection during message propagation by adding or cutting edges, or by assigning different weights to each edge [9, 20, 34]. In addition, several methods employ advanced strategies to address the homophily discrepancy in message passing, such as gradient-based filtering [12], meta-learning [37, 7], and data augmentation [28, 45]. These methods overlook the homophily differences among the labeled normal nodes in semi-supervised settings. our proposed RHO is the first work to reveal this issue and address it by a robust homophily learning approach.
@@ 3 Problem Statement
Notations. Given an attributed graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{X})$, where $\mathcal{V}$ denotes the node set with $v_{i}\in\mathcal{V}$ and $\left|\mathcal{V}\right|=N$, $\mathcal{E}$ denotes the edge set, and $\mathbf{X}=\left[\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{N}\right]\in\mathbb{R}^{N\times M}$ is a set of node attributes. Each node $v_{i}$ has a $M$-dimensional feature representation $\mathbf{x}_{i}$. The topological structure of $\mathcal{G}$ is represented by an adjacency matrix $\mathbf{A}$. $\mathbf{D}$ denotes a degree matrix which is a diagonal matrix with $\mathbf{D}_{ii}=\sum_{j}\mathbf{A}_{ij}$. Normalized Laplacian matrix $\mathbf{L}$ is defined by $\mathbf{L}=\mathbf{I}_{N}-\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}$, where $\mathbf{I}_{N}\in\mathbb{R}^{N\times N}$ denotes an identity matrix. $\hat{\mathbf{A}}$ is the normalized adjacency matrix. The corresponding degree matrix $\hat{\mathbf{D}}$ and Laplacian matrix $\hat{\mathbf{L}}$ are defined as $\hat{\mathbf{D}}=\mathbf{D}+\mathbf{I}_{N}$ and $\hat{\mathbf{L}}=\mathbf{I}_{N}-{\hat{\mathbf{D}}}^{-1/2}\hat{\mathbf{A}}{\hat{\mathbf{D}}}^{-1/2}$, respectively. The homophily ratio $\mathcal{H}$ quantifies the homophily level in a graph, defined as the fraction of edges connecting nodes with the same class label, i.e., =|$\mathcal{H}=\frac{\left|\left\{(v_{i},v_{j}):(v_{j},v_{i})\in\mathcal{E}\wedge y_{i}=y_{j}\right\}\right|}{|\mathcal{E}|}$. A graph exhibits high homophily when $\mathcal{H}\rightarrow 1$, whereas low homophily (or high heterophily) corresponds to $\mathcal{H}\rightarrow 0$.
@@
Graph Filter. The Laplacian matrix $\hat{\mathbf{L}}$ can be decomposed as $\hat{\mathbf{L}}=\mathbf{U}\Lambda\mathbf{U}^{\top}$, where $\mathbf{U}=(\mathbf{u}_{1},\mathbf{u}_{2},\ldots,\mathbf{u}_{N})$ is a complete set of orthonormal eigenvectors known as graph Fourier modes and $\Lambda=\operatorname{diag}\left(\left\{\lambda_{i}\right\}_{i=1}^{N}\right)$ is a diagonal matrix of the eigenvalues of $\hat{\mathbf{L}}$, where $\lambda_{1}\leq\lambda_{2}\leq\cdots\leq\lambda_{N}$ and $\lambda_{i}\in[0,1.5]$ [42]. In graph signal processing (GSP), $\{\lambda_{i}\}$ and $\{\mathbf{u}_{i}\}$ are called frequencies and frequency components of graph $\mathcal{G}$. Taking the eigenvectors of normalized Laplacian matrix as a set of bases, graph Fourier transform of a signal $\mathbf{x}\in\mathbb{R}^{N}$ on graph $\mathcal{G}$ is defined as $\hat{\mathbf{x}}=\{\hat{x}_{1},\cdots,\hat{x}_{N}\}=\mathbf{U}^{T}\mathbf{x}$, and the inverse graph Fourier transform is $\mathbf{x}=\mathbf{U}\hat{\mathbf{x}}$. The convolution between the signal $\mathbf{x}$ and convolution kernel $g$ is as follows:
$g*\mathbf{x}=\mathbf{U}\left(\left(\mathbf{U}^{T}g\right)\odot\left(\mathbf{U}^{T}\mathbf{x}\right)\right)=\mathbf{U}g_{\theta}\mathbf{U}^{T}\mathbf{x},$ (1)
where $\odot$ is an element-wise product and $g_{\theta}$ is a convolutional kernel in the spectral domain.
@@ 4.1 Overview of the Proposed Approach RHO
As shown in Fig. 2, RHO consists of three main components: (1) Adaptive frequency response (AdaFreq) filters for heterogeneous normal pattern learning, which is simultaneously applies on both cross-channel and channel-wise view; (2) Graph normality alignment (GNA) for aligning the learned normality from two views using a contrastive learning objective; and (3) the model is lastly optimized with a widely-used one-class objective, along with a contrastive loss in GNA, to learn robust homophily patterns on GAD datasets with diverse levels of homophily.
@@ 4.2 AdaFreq: Adaptive Frequency-response Filters
Let $\{\lambda_{m}\}$ and $\{\mathbf{u}_{m}\}$ be the graph frequencies and frequency components respectively, $\beta_{m}$ is the projection coefficient of signal $\mathbf{x}$ onto the $m$-th eigenvector $\mathbf{u}_{m}$, assuming $\beta_{m}>0$, then we have $g\left(\lambda_{m}\right)=\frac{\sum_{i\in\mathcal{V}_{l}}u_{m}(i)}{\beta_{m}\sum_{i\in\mathcal{V}_{l}}u_{m}(i)^{2}}$ for the filter $g\left(\lambda\right)$, indicating that frequencies where normal nodes show coherent spectral behavior (i.e., $u_{m}(i)$ values agree in sign/magnitude) are amplified, while inconsistent frequency components are suppressed.
The proof can be found in App. A. According to Theorem 1, the designed adaptive filter $g\left(\lambda\right)$ can preserve the frequency components with consistent spectral behavior of labeled normal nodes by training the parameter $k$. When $k>0$, $g(\lambda)$ decreases with respect to $\lambda$, meaning that frequency components associated with larger eigenvalues (i.e., high-frequency components) are suppressed, while low-frequency components are preserved. Conversely, when $k<0$, $g(\lambda)$ emphasizes high-frequency components. The magnitude of $k$ controls the strength of frequency responses. When $k=0$, $g(\lambda)$ retains raw frequency signals. When $K$ layers are stacked, $g(\lambda)=\prod_{i=1}^{K}(1-k_{i}\lambda)$, a set of trainable parameters $\{k_{1},k_{2},\cdots,k_{K}\}$ can produce more complex frequency responses. Therefore, $g(\lambda)$ can dynamically learn the parameter $k$ to capturing the consistent frequency components of normal nodes with diverse homophily, enabling robust homophily representation learning for GAD.
Figure 3: Anomaly score distributions and t-SNE visualizations of node embeddings generated by the cross-channel view ((a) and (d)), the channel-wise view ((b) and (e)), and the full RHO model ((c) and (f)) on Amazon [9].
@@
Homophily Learning across the Channels. According to Eq. (1) and (2), AdaFreq can be utlized to learn a single learnable parameter $k\in\mathbb{R}$ shared across all node attributes in the cross-channel view. The learned cross-channel representations, denoted by $\mathbf{H}_{ccr}^{(t)}$ for the $t$-th layer, can be defined as:
@@
$\mathcal{L}_{ccr}=\frac{1}{|\mathcal{V}_{l}|}\sum_{i\in\mathcal{V}_{l}}\left\|\mathbf{h}_{ccr}^{(T)}(i)-\bm{c}_{ccr}\right\|^{2}+\sum_{t=1}^{T}||W^{(t)}_{ccr}||_{F}^{2},$ (4)
@@
where $[\mathbf{h}_{1}^{(t-1)},\mathbf{h}_{2}^{(t-1)},\cdots,\mathbf{h}_{d}^{(t-1)}]=\mathbf{H}_{cwr}^{(t-1)}\in\mathbb{R}^{N\times d}$, and $\mathbf{H}_{cwr}^{(0)}=f_{\theta}(\mathbf{X})$.
@@
$\mathcal{L}_{cwr}=\frac{1}{|\mathcal{V}_{l}|}\sum_{i\in\mathcal{V}_{l}}\left\|\mathbf{h}_{cwr}^{(T)}(i)-\bm{c}_{cwr}\right\|^{2}++\sum_{t=1}^{T}||W^{(t)}_{cwr}||_{F}^{2},$ (7)

### v2

@@ Abstract
Current semi-supervised graph anomaly detection (GAD) methods utilizes a small set of labeled normal nodes to identify abnormal nodes from a large set of unlabeled nodes in a graph. These methods posit that 1) normal nodes share a similar level of homophily and 2) the labeled normal nodes can well represent the homophily patterns in the entire normal class. However, this assumption often does not hold well since normal nodes in a graph can exhibit diverse homophily in real-world GAD datasets. In this paper, we propose RHO, namely Robust Homophily Learning, to adaptively learn such homophily patterns. RHO consists of two novel modules, adaptive frequency response filters (AdaFreq) and graph normality alignment (GNA). AdaFreq learns a set of adaptive spectral filters that capture different frequency components of the labeled normal nodes with varying homophily in the channel-wise and cross-channel views of node attributes. GNA is introduced to enforce consistency between the channel-wise and cross-channel homophily representations to robustify the normality learned by the filters in the two views. Experiments on eight real-world GAD datasets show that RHO can effectively learn varying, often under-represented, homophily in the small labeled node set and substantially outperforms state-of-the-art competing methods. Code is available at https://github.com/mala-lab/RHO.
@@ 1 Introduction
Graph anomaly detection (GAD) has received increasing attention in recent years due to its wide range of applications, such as fraud detection in finance, review spam detection, and abusive user detection [akoglu2015graph, ma2021comprehensive, liu2022bond, qiao2024deep]. Semi-supervised GAD, which aims to leverage a small set of labeled normal nodes to identify abnormal nodes from a large set of unlabeled nodes in a graph, is among the most realistic problem settings in real-life applications [pang2021deep, qiao2024generative, qiao2024deep]. This is because GAD datasets are typically dominated by normal nodes, which can greatly facilitate the annotation of a few normal nodes, e.g., if we randomly sample a few nodes from a large graph, most nodes would be normal. Current semi-supervised GAD methods are generally built upon two key assumptions: (1) all normal nodes exhibit a similar level of homophily, and (2) the labeled normal nodes can well represent the overall homophily pattern of the entire graph [fan2020anomalydae, ding2019deep, qiao2024generative, ding2021inductive, wang2021one]. However, these two assumptions often do not hold well in the real-world GAD datasets. This is because although normal nodes generally exhibit high homophily, there can be large variations of homophily across the normal nodes, some of which can possess relatively low homophily (see Fig. 1a for visualization of this observation on the Amazon [dou2020enhancing] and Elliptic [weber2019anti]. Additional visualizations can be found in App. D.1). As a result, the GAD methods that rely on the aforementioned assumptions learn inaccurate homophily patterns of the normal class, leading to the misclassificaiton of low-homophily normal nodes as abnormal. This is particularly true when they use the popular neighborhood aggregation mechanisms that are prone to produce oversmooth homophily representations [qiao2024deep].
@@
Figure 1: (a) Homophily distribution of normal nodes in Amazon [dou2020enhancing] and Elliptic [weber2019anti]. (b) and (c) AUC comparison of GCN filter [wang2021one], BWGNN filter [tang2022rethinking], and AdaFreq filter on the two datasets under three cases of homophily levels. The three cases are normal nodes sampled from the sets of (low-homophily, high-homophily) normal nodes, where nodes with homophily greater than 0.9 for Amazon and 0.7 for Elliptic are considered as high homophily. The datasets in the three cases are sampled using a ratio of (80%, 20%), (50%, 50%), and (20%, 80%) to the two node sets, respectively.
RHO consists of two novel modules, including adaptive freqency response filters (AdaFreq) and graph normality alignment (GNA). Unlike existing filters that are bounded by prior knowledge of a specific homophily, AdaFreq learns a set of adaptive filters with learnable parameters to robustly fit normal nodes of varying levels of homophily. For example, as illustrated in Fig. 1b and 1c, conventional GCN filters [wang2021one] rely on a low-frequency assumption of normal nodes, while the spectral filters in BWGNN [tang2022rethinking] learns a pre-defined type of frequency response, both of which fail to learn generalized representations for normal nodes across three different homophily variations; AdaFreq is instead optimized with learnable parameters to fit adaptively to the normal nodes with three different homophily cases, showing much more robustness than the other two methods. In RHO, AdaFreq is utilized to learn such adaptive filters in both channel-wise and cross-channel views of node attributes, enabling the learning of heterogeneous normal patterns in diverse feature channels of the labeled normal nodes.
The set of filters learned in AdaFreq can differ from each other substantially. To obtain robust yet consistent representations of the normal nodes, the GNA module is introduced to enforce consistency of homophily representations learned across the filters. To this end, GNA constructs positive pairs using the representations from the respective channel-wise and cross-channel views at the corresponding nodes, and then maximizes the similarity between the representations learned from these two views while minimizing the similarity of negative pairs composed by representations from different nodes.
In doing so, RHO learns heterogeneous normal representations of normal nodes with varying degrees of homophily via AdaFreq, while ensuring the consistency of the learned normality via GNA. In summary, this work makes the following three main contributions.
@@ 2.1 Graph Anomaly Detection (GAD)
Non-spectral GAD Methods. Many non-sepctral GAD methods apply traditional anomaly detection techniques on a GNN backbone, such as reconstruction [ding2019deep, fan2020anomalydae], adversarial learning [ding2021inductive, chen2020generative], one-class classification [zhou2021subtractive, wang2021one, cai2024deep]. Other methods are based on measures designed for GAD, such as the recently proposed affinity-based methods [pan2023prem, hezhe2023truncated, pan2025label, niu2024zero]. These methods are typically designed for the unsupervised setting, where no labeled nodes are available. They can be easily adapted to the semi-supervised setting by refining their objective to the labeled normal nodes. GGAD [qiao2024generative] is the first method specifically designed for the semi-supervised setting by generating the outliers and training a discriminative one-class classifier with the given labeled normal nodes. These methods failed to consider the homophily discrepancy among the normal nodes. Our proposed RHO addresses this by employing an adaptive filter that learns normal representations from both cross-channel and channel-wise views through a one-class optimization objective.
Spectral GAD Methods. The spectral GAD methods focus on the filter design in GNNs to have a more effective learning the representations of nodes [chai2022can, tang2023gadbench, gao2023addressing]. Among them, AMNet [chai2022can], BWGNN [tang2023gadbench], and GHRN [gao2023addressing] aim to utilize the frequency signal using different spectral-based GNNs to distinguish between normal and abnormal nodes. Although these spectral-based GAD methods have achieved remarkable success [liu2025pre, dong2025spacegnn], they are typically designed for fully supervised setting where both labeled normal and abnormal nodes are available during training. On the other hand, the filters they utilize learn only predefined frequency information, making them difficult to learn expressive representations for normal nodes of different frequency components. RHO leverages AdaFreq to assign learnable parameters on different feature channels to effectively learn heterogeneous normal representations from the set of normal nodes of varying homophily.
@@ 2.2 Graph Homophily Modeling for GAD
Graph homophily modeling methods aim to address the homophily discrepancy in graphs, where target nodes often connect to nodes from different classes [zheng2022graph, gong2024survey, luan2024heterophily]. This inter-class connectivity negatively affects the quality of node representations within each class. The problem becomes more pronounced in GAD, where the homophily gap is typically large due to the inherent imbalance in GAD datasets [zheng2022graph, luan2024heterophilic, qiao2024deep]. Existing studies on Graph homophily modeling for GAD primarily focus on neighbor selection during message propagation by adding or cutting edges, or by assigning different weights to each edge [dou2020enhancing, liu2021pick, hezhe2023truncated]. In addition, several methods employ advanced strategies to address the homophily discrepancy in message passing, such as gradient-based filtering [gao2023addressing], meta-learning [qin2022explainable, dong2022bi], and data augmentation [chen2024consistency, meng2023generative, zhou2023improving]. These methods overlook the homophily differences among the labeled normal nodes in semi-supervised settings. our proposed RHO is the first work to reveal this issue and address it by a robust homophily learning approach.
@@ 3 Problem Statement
Notations. Given an attributed graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{X})$, where $\mathcal{V}$ denotes the node set with $v_{i}\in\mathcal{V}$ and $\left|\mathcal{V}\right|=N$, $\mathcal{E}$ denotes the edge set, and $\mathbf{X}=\left[\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{N}\right]\in\mathbb{R}^{N\times M}$ is a set of node attributes. Each node $v_{i}$ has a $M$-dimensional feature representation $\mathbf{x}_{i}$. The topological structure of $\mathcal{G}$ is represented by an adjacency matrix $\mathbf{A}$. $\mathbf{D}$ denotes a degree matrix which is a diagonal matrix with $\mathbf{D}_{ii}=\sum_{j}\mathbf{A}_{ij}$. Normalized Laplacian matrix $\mathbf{L}$ is defined by $\mathbf{L}=\mathbf{I}_{N}-\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}$, where $\mathbf{I}_{N}\in\mathbb{R}^{N\times N}$ denotes an identity matrix. $\hat{\mathbf{A}}$ is the normalized adjacency matrix. The corresponding degree matrix $\hat{\mathbf{D}}$ and Laplacian matrix $\hat{\mathbf{L}}$ are defined as $\hat{\mathbf{D}}=\mathbf{D}+\mathbf{I}_{N}$ and $\hat{\mathbf{L}}=\mathbf{I}_{N}-{\hat{\mathbf{D}}}^{-1/2}\hat{\mathbf{A}}{\hat{\mathbf{D}}}^{-1/2}$, respectively. The node homophily [pei2020geom] for a node $v$ defined as $\mathcal{H}_{v}=\frac{\left|\left\{u\mid u\in\mathcal{N}_{v},y_{u}=y_{v}\right\}\right|}{d_{v}}$, where $d_{v}$ is the number of neighbors of node $v$, $\mathcal{N}_{v}$ is the set of adjacent nodes of $v$, and $y_{v}$ represents the label of node $v$.
@@
Graph Filter. The Laplacian matrix $\hat{\mathbf{L}}$ can be decomposed as $\hat{\mathbf{L}}=\mathbf{U}\Lambda\mathbf{U}^{\top}$, where $\mathbf{U}=(\mathbf{u}_{1},\mathbf{u}_{2},\ldots,\mathbf{u}_{N})$ is a complete set of orthonormal eigenvectors known as graph Fourier modes and $\Lambda=\operatorname{diag}\left(\left\{\lambda_{i}\right\}_{i=1}^{N}\right)$ is a diagonal matrix of the eigenvalues of $\hat{\mathbf{L}}$, where $\lambda_{1}\leq\lambda_{2}\leq\cdots\leq\lambda_{N}$ and $\lambda_{i}\in[0,1.5]$ [wu2019simplifying]. Taking the eigenvectors of normalized Laplacian matrix as a set of bases, graph Fourier transform of a signal $\mathbf{x}\in\mathbb{R}^{N}$ on graph $\mathcal{G}$ is defined as $\hat{\mathbf{x}}=\{\hat{x}_{1},\cdots,\hat{x}_{N}\}=\mathbf{U}^{T}\mathbf{x}$, and the inverse graph Fourier transform is $\mathbf{x}=\mathbf{U}\hat{\mathbf{x}}$. The convolution between the signal $\mathbf{x}$ and convolution kernel $f$ is as follows:
$f*\mathbf{x}=\mathbf{U}\left(\left(\mathbf{U}^{T}f\right)\odot\left(\mathbf{U}^{T}\mathbf{x}\right)\right)=\mathbf{U}g_{\theta}\mathbf{U}^{T}\mathbf{x},$ (1)
where $\odot$ is an element-wise product and $g_{\theta}$ is a diagonal matrix representing the convolution kernel in the spectral domain, serving as a substitute of $\mathbf{U}^{T}f$.
@@ 4.1 Overview of the Proposed Approach RHO
As shown in Fig. 2, RHO consists of three main components: (1) Adaptive frequency response (AdaFreq) filters for heterogeneous normal pattern learning, which is simultaneously applies to both cross-channel and channel-wise view; (2) Graph normality alignment (GNA) for aligning the learned normality from two views using a contrastive learning objective; and (3) the model is lastly optimized with a widely-used one-class objective, along with a contrastive loss in GNA, to learn robust homophily patterns on GAD datasets with diverse levels of homophily. We also summarize the workflow of RHO and provide a detailed algorithmic description in App. E.
@@ 4.2 AdaFreq: Adaptive Frequency-response Filters
Let $\{\lambda_{m}\}$ and $\{\mathbf{u}_{m}\}$ be the graph frequencies and frequency components respectively, $\beta_{m}$ is the projection coefficient of signal $\mathbf{x}$ onto the $m$-th eigenvector $\mathbf{u}_{m}$, then we have $g\left(\lambda_{m}\right)=\frac{\sum_{i\in\mathcal{V}_{l}}u_{m}(i)}{\beta_{m}\sum_{i\in\mathcal{V}_{l}}u_{m}(i)^{2}}$ for the filter $g\left(\lambda\right)$, indicating that frequencies where normal nodes show coherent spectral behavior (i.e., $u_{m}(i)$ values agree in sign/magnitude) are amplified, while inconsistent frequency components are suppressed.
The proof can be found in App. A. According to Theorem 1, the designed adaptive filter $g\left(\lambda\right)$ can preserve the frequency components with consistent spectral behavior of labeled normal nodes by training the parameter $k$. When $k>0$, $g(\lambda)$ decreases with respect to $\lambda$, meaning that frequency components associated with larger eigenvalues (i.e., high-frequency components) are suppressed, while low-frequency components are preserved. Conversely, when $k<0$, $g(\lambda)$ emphasizes high-frequency components. The magnitude of $k$ controls the strength of frequency responses. When $k=0$, the filter $g(\lambda)=1$ acts as an all-pass filter preserving the raw graph signal. When $K$ layers are stacked, $g(\lambda)=\prod_{i=1}^{K}(1-k_{i}\lambda)$, a set of trainable parameters $\{k_{1},k_{2},\cdots,k_{K}\}$ can produce more complex frequency responses. Therefore, $g(\lambda)$ can dynamically learn the parameter $k$ to capturing the consistent frequency components of normal nodes with diverse homophily, enabling robust homophily representation learning for GAD.
Figure 3: Anomaly score distributions and t-SNE visualizations of node embeddings generated by the cross-channel view ((a) and (d)), the channel-wise view ((b) and (e)), and the full RHO model ((c) and (f)) on Amazon [dou2020enhancing].
@@
Homophily Learning across the Channels. According to Eq. (1) and (2), AdaFreq can be utilized to learn a single learnable parameter $k\in\mathbb{R}$ shared across all node attributes in the cross-channel view. The learned cross-channel representations, denoted by $\mathbf{H}_{ccr}^{(t)}$ for the $t$-th layer, can be defined as:
@@
[…]
@@
where $\mathbf{h}_{1}^{(t-1)},\mathbf{h}_{2}^{(t-1)},\cdots,\mathbf{h}_{d}^{(t-1)}]=\mathbf{H}_{cwr}^{(t-1)}\in\mathbb{R}^{N\times d}$, and $\mathbf{H}_{cwr}^{(0)}=f_{\theta}(\mathbf{X})$.
@@
[…]